OpenStudy (anonymous):

Which of the following is a polynomial with roots – square root of 3, square root of 3, and –2 ? x^3 – 2x^2 – 3x + 6 x3 – 3x^2 – 5x + 15 x^3 + 2x2 – 3x – 6 x^3 + 3x^2 – 5x – 15

OpenStudy (anonymous):

You can do this 2 ways. Do you have any idea how?

OpenStudy (anonymous):

\[-\sqrt{3}, \sqrt{3} and -2\]

OpenStudy (anonymous):

Not really andras

OpenStudy (anonymous):

You can check each individually but this would take forever. (some time) Or you are wise and know that if a polynomial has three roots: lets say a,b,c than its equation is (x-a)(x-b)(x-c) Do you see how this can help you?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

I don't see the a b and c values :l

OpenStudy (anonymous):

\[(x--\sqrt{3})(x-\sqrt{3})(x--2)=solution \]

OpenStudy (anonymous):

ok so whats our first step?

OpenStudy (anonymous):

This was the first step. The second step is the expand those brackets!

OpenStudy (anonymous):

As the solutions are given it is easy to see that it can only be x^3 + 2x^2 – 3x – 6 I only taken a look at -6. The only part of this polynomial without x is the: \[(--\sqrt{3})(-\sqrt{3})(--2)=-6\]

OpenStudy (anonymous):

Thanks man :)